Solve by setting up the proper equation to describe the facts given and then carrying out the mathematical calculations to solve for the unknown variable(s).
The sum of two numbers is 36. Their difference is 4. Determine the two unknown numbers.
(NOTE: Show answer in the following format # and #. If not entered in this format, will be scored incorrect.)

Found 2 solutions by mananth, jim_thompson5910:
Answer by mananth(16946)

x= one number
y= second number

1 x + 1 y = 36 .............1

1 x -1 y = 4 .............2
Eliminate y
Add the two equations
2 x = 40
/ 2
x = 20
plug value of x in (1)
1 x + 1 y = 36
20 + y = 36
y = 36 -20
y = 16
y = 16 1
x= 20 one number
y= 16 second number
m.ananth@hotmail.ca

Answer by jim_thompson5910(35256)

Hint: the system of equations is

x+y = 36
x-y = 4

Let me know if you need more help.


Feel free to email me at jim_thompson5910@hotmail.com
or you can visit my website here: http://www.freewebs.com/jimthompson5910/home.html

Thanks,

Jim